OpenStudy (anonymous):

I'm in need of help.... Solve using two different methods. Explain which method you found to be more efficient: A. 3x - 9y = 3 6x - 3y = -24 B. 7x - 3y = 20 5x + 3y = 16 C. y = (1/2)x - 6 2x + 6y = 19

OpenStudy (anonymous):

Usually I'm good with math, but this year I'm getting creamed...

OpenStudy (studygurl14):

the two methods are elimination and substitution. Do both per question, then say which one is easier

OpenStudy (studygurl14):

Do you know how to do elimination and substtuion?

OpenStudy (anonymous):

Not really... learned this stuff last week and got a concussion at work... head throbbing to much to remember anything. /:

OpenStudy (anonymous):

lol i read it wrong didn't I? it's not 3 equations and 3 multiple choice answers...

OpenStudy (anonymous):

Still you can use that as a template for each equation. Basically \[ax+by=cz\] and \[dx+ey=fz\] then \[x=\frac{ cz-by }{ a }\] can be put into \[d(\frac{ cz-by }{ a }) +ey = fz\]

OpenStudy (anonymous):

That is substitution

OpenStudy (anonymous):

okay... this'll be a long day ha

OpenStudy (anonymous):

You could always just use that formula I created, \[6(\frac{ 3+9y }{ 3 } )-9y=3\]

OpenStudy (anonymous):

That's making a lot more sense haha... sorry, like I said, my mind is boggled and head hurts from the "work place accident"

OpenStudy (anonymous):

So, what about the 3x - 9y part?

OpenStudy (studygurl14):

A. 3x - 9y = 3 6x - 3y = -24 ELIMNINATION:  SUBSTITUTION:  This is the first one. What is easier to you -- elimination or subtraction?